Feemster Corporation manufactures and sells a single product. The company uses units as the measure of activity in its budgets and performance reports. During October, the company budgeted for 5,910 units, but its actual level of activity was 5,860 units. The company has provided the following data concerning the formulas used in its budgeting and its actual results for October:

Data used in budgeting:

Fixed element per month Variable element per unit
Revenue - $ 30.00
Direct labor $ 0 $ 5.70
Direct materials 0 8.70
Manufacturing overhead 37,100 1.40
Selling and administrative expenses 27,100 .90
Total expenses $ 64,200 $ 16.70

Actual results for October:

Revenue $ 176,595
Direct labor $ 32,010
Direct materials $ 50,300
Manufacturing overhead $ 46,275
Selling and administrative expenses $ 32,170

The activity variance for selling and administrative expenses in October would be closest to:

Multiple Choice

  • $45 U

  • $249 F

  • $249 U

  • $45 F
Answer #1
Activity variance for selling & administrative expenses
selling & administrative exp (planning budget) = (0.90*5910) +27100 = 32419
selling & administrative exp (flexible budget) = (0.90*5860)+27100 = 32374
activity variance =selling & administrative exp (planning budget)-Selling & adm exp (flexible budget)
activity variance = 32419-32374 = 45 F
Option 45 F is CORRECT

Please give a thumbs up if it is helpful & let me know if any doubt